# A two-phase model of galaxy formation: III. The formation of globular clusters

### v2

@@ 5 Summary and discussion
    A number of observational results of GCs in the local Universe are reproduced, and more results are predicted and to be tested by future observations. These include the mass and age distributions of GCs in different galaxies; the size-mass relation of GCs; the relation between the total mass in GCs ($M_{*}^{\text{(GCs)}}$) and the host halo/galaxy mass; the bimodal metallicity distribution; the fraction of Pop-I/II GCs as a function halo mass; and the spatial profile of GCs within host galaxies (§4).
@@
    The linear relation between $M_{*}^{\text{(GCs)}}$ and the host halo mass is set up by the metal enrichment process, and is preserved by the central limit theorem (§4.4).
@@
A number of assumptions in the model of GC have to be verified by observations and/or hydrodynamical simulations. For example, the independent sampling of sub-cloud density and mass (§3.1) is designed to mimic the mass-independent density of virialized dark matter halos given by the spherical collapse model, while dissipative processes in SGCs may introduce a density-mass correlation. This can be checked by hydrodynamical simulation with a resolution that is sufficient for resolving the formation of sub-clouds. Another example is the role of radiative and wind feedback from the internal star formation of a sub-cloud. We argued that such feedback produces the flat size-mass relation of GCs (§3.4). These feedback channels have been simulated in hydrodynamic simulations (e.g. Hopkins et al., 2020; Li et al., 2020), and can be tested by measuring the correlation of gas clouds and young star clusters in observations.
@@ Appendix B Calibration of the Model
The most uncertain part of the model is the process of metal enrichment, partly due to the limited constraints from high-redshift observations (e.g. Curti et al., 2020; Sanders et al., 2021; Curti et al., 2023; Venturi et al., 2024), and partly due to the potential systematic differences in metallicity in various regions of the host galaxy (Wang et al., 2019, 2022; Venturi et al., 2024), particularly within the dense sub-clouds. Given the lack of constraints, we propose a heuristic model in §2.4, motivated by the gas-regulator scenario, and adapt it to account for conditions prevalent at high redshifts. The resulting MZR of this model is required to reproduce observed trends, including the slow redshift evolution at $z\gtrsim 4$ (e.g. Nakajima et al., 2023), the slope of 1/3absent13$\approx$ 1/3≈ 1 / $32/$ 3 at the low-stellar-mass end (see e.g. figure 3 of Curti et al., 2020), and a bending at the high-stellar-mass end (e.g. Tremonti et al., 2004; Mannucci et al., 2010). All parameters involved in the metal enrichment process affects GC formation via the Pop-II channel (see §3.5), and we adjust them within acceptable ranges that meet the aforementioned requirements to reproduce the GC-to-halo mass ratio and the Pop-I fraction. For example, the pink curves in Fig. 21 show the results of a model variant, denoted as ‘Weak Confinement’, with an escape velocity of $v_{\rm esc}=250\,{\rm{km\,s^{-1}}}$, which is larger than the fiducial value of $75\,{\rm{km\,s^{-1}}}$. This variant tends to produce a more metal-poor environment, therefore enhancing the formation of Pop-II GCs. Consequently, the changes in the GC-to-halo mass ratio and Pop-I fraction are both significant, resulting in a clear inconsistency with observations.

### v3

@@ 5 Summary and discussion
    With calibration by observations, a number of observational results of GCs in the local Universe are reproduced, and more results are predicted and to be tested by future observations. The mass and age distributions of GCs produced by the model cover those observed in MW and M31, but do not reproduce those seen in M51, an interacting galaxy. The size-mass relation of GCs, with its normalization tuned to match observations, can be explained as a result of gravity-feedback balance in the star formation of sub-clouds, provided that subsequent dynamical evolution does not drive GCs off the birth-time relation. The relation between the total mass in GCs ($M_{*}^{\text{(GCs)}}$) and the host halo (galaxy) mass is predicted to be linear (non-linear). The bimodal distribution in GC metallicity is produced naturally by the two-channel formation, and the fraction of Pop-I/II GCs as a function of halo mass reflects the enrichment history of the host galaxy. The predicted spatial profile of GCs within the host galaxy is shaped by both in-situ and ex-situ, and both fast-phase and slow-phase, formation of the host galaxy. The modeled profiles match those of MW and M31, but show discrepancies with those observed for dwarfs and BCGs (§4).
@@
    The linear relation between $M_{*}^{\text{(GCs)}}$ and the host halo mass is set by the metal enrichment process, and is preserved by the central limit theorem (§4.4).
@@
A number of assumptions in the model of GC have to be verified or extended by observations and/or hydrodynamical simulations. For example, the independent sampling of sub-cloud density and mass (§3.1) is designed to mimic the mass-independent density of virialized dark matter halos given by the spherical collapse model, while dissipative processes in SGCs may introduce a density-mass correlation. Meanwhile, the sub-cloud mass function can be modified by the processes discussed in §3.1, and may thus depend on the environment. These can be explored by hydrodynamical simulation with a resolution that is sufficient for resolving the formation of sub-clouds. Another example is the role of radiative and wind feedback from the internal star formation of a sub-cloud. We argued that such feedback produces the flat size-mass relation of GCs (§3.4). These feedback channels have been simulated in hydrodynamic simulations (e.g. Hopkins et al., 2020; Li et al., 2020), and can be tested by measuring the correlation of gas clouds and young star clusters in observations. The final example is the spatial distribution of GCs within their host galaxies, which is modeled semi-empirically with three-components described in §4.6 and should be improved by directly tracing individual GCs within realistic gravitational potential. The comparison of the spatial distribution with observations can be made by forward modeling that mimics the selection and projection in observations.
@@ Appendix B Calibration of the Model
The most uncertain part of the model is the process of metal enrichment, partly due to the limited constraints from high-redshift observations (e.g. Curti et al., 2020; Sanders et al., 2021; Curti et al., 2024; Venturi et al., 2024), and partly due to the potential systematic differences in metallicity in various regions of the host galaxy (Wang et al., 2019, 2022; Venturi et al., 2024), particularly within the dense sub-clouds. Given the lack of constraints, we propose a heuristic model in §2.4, motivated by the gas-regulator scenario, and adapt it to account for conditions prevalent at high redshifts. The resulting MZR of this model is required to reproduce observed trends, including the slow redshift evolution at $z\gtrsim 4$ (e.g. Nakajima et al., 2023), the slope of 1/3absent13$\approx$ 1/3≈ 1 / $32/$ 3 at the low-stellar-mass end (see e.g. figure 3 of Curti et al., 2020), and a bending at the high-stellar-mass end (e.g. Tremonti et al., 2004; Mannucci et al., 2010). All parameters involved in the metal enrichment process affects GC formation via the Pop-II channel (see §3.5), and we adjust them within acceptable ranges that meet the aforementioned requirements to reproduce the GC-to-halo mass ratio and the Pop-I fraction. For example, the pink curves in Fig. 21 show the results of a model variant, denoted as ‘Weak Confinement’, with an escape velocity of $v_{\rm esc}=250\,{\rm{km\,s^{-1}}}$, which is larger than the fiducial value of $75\,{\rm{km\,s^{-1}}}$. This variant tends to produce a more metal-poor environment, therefore enhancing the formation of Pop-II GCs. Consequently, the changes in the GC-to-halo mass ratio and Pop-I fraction are both significant, resulting in a clear inconsistency with observations.